Fermi National Laboratory
 
Operation Summary for April 6, 2012 - April 9, 2012
 

NuMI reported receiving 8.61E18 protons on target for the period from 4/2/12 to 4/9/12.

Friday April 6

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, MiniBooNE, and SeaQuest experiments taking beam.

During the day shift, Operations established beam to the Meson FTBF T-992 experiment at 9:17 AM. At 9:59 AM, a Linac Klystron RF station (KRF4) tripped off due to high tank humidity; it reset. Muon Ring personnel began a study at 10:21 AM. KRF4 tripped off again on tank humidity. Linac personnel reported at 10:34 AM that KRF4 would have to have its tube replaced; Operations turned off Linac beam at 11:13 AM. Operators prepared MI and SeaQuest for access at 11:16 AM. Operators prepared the Muon Ring for access at 11:17 AM. Operators prepared Booster for an access at 12:28 PM to replace a Power Amplifier for BRF9. Power Supply personnel began working on NuMI power supply E:V118 at 1:01 PM. Booster came out of access at 3:04 PM.

During the evening shift, Water Group personnel reported at 4:24 PM that they had fixed the MI RF LCW leaks. Operators began ramping the MI power supplies at 5:23 PM. A Linac person reported at 5:59 PM that they had completed their work on KRF4. Operations established beam to NuMI at 6:51 PM and to SeaQuest at 6:56 PM. Operations reported that towards the end of the shift KRF7 began tripping off on reflected power; Operators reset the station’s water temperature loop. Operations disabled beam to the T-992 experiment.

Saturday April 7

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, and SeaQuest.

During the midnight shift, Operations reported beam problems in the 400 MeV area at 12:20 AM; a Linac person believed the problem was with KRF4 and he and several other Linac personnel came in to work on the problem. At 3:23 AM, the NuMI beam permit tripped off due to a power supply’s (E:V118) current drifting low; everything reset. Operations had beam establish to MiniBooNE, NuMI, and SeaQuest by 3:45 AM.

During the day shift, Operations established beam to T-992 at 12:17 PM. Operators completed a Booster injection tune at 3:19 PM. T-992 made an access during this shift.

During the evening shift, at 5:09 PM, Operators reported a problem with the I- Source extractor current slowly rising; a Linac person asked Operators to make an adjustment.

Sunday April 8

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-992.

During the midnight shift, at 12:28 AM, KRF4 tripped off for the first of two times in about 15 minutes; it reset. At 12:29 AM, the NuMI beam permit tripped off for the first of two times due to V118 in six and a half hours; it reset. Operations disabled beam to T-992 at 2:50 AM. Several MIRF and Booster RF trips occurred during this shift.

During the day shift, at 10:44 AM, the NuMI beam permit tripped off for the first of three times due V118 in about 18 minutes; it reset. Operators reported that the H- Source (the backup source) high voltage had been tripping off; it reset, but continued to be unstable.

During the evening shift, Operators reported at 8:41 PM that they had added water to several RF stations. At 9:35 PM, the NuMI beam permit tripped off for the first of two times in about two hours due to V118; it reset.

Monday April 9

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, and SeaQuest.

During the midnight shift, at 6:02 AM, Operations reported that there had been 54 trips of the MiniBooNE beam permit due to what Operators think was a Booster kicker (MP02) problem. At 7:29 AM, the NuMI beam permit tripped off due to V118; it reset.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo agrees with the plans of Muon Ring and MI personnel to conduct studies. SeaQuest continues to commission their beam line. MTA personnel will want beam this week.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/09/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory